# Conditional Video Generation for High-Efficiency Video Compression

### v3

@@ 3. Conditional Frame Generation at Decoder
Fig. 5 provides a qualitative comparison of video compression performance across different methods, specifically between Ours, H.264, and the Ground Truth. The selected samples illustrate how our method preserves semantic and structural integrity under four complex scenarios.
(a) Scene: A Young Man on a Cliff
In this natural outdoor scene, the Ground Truth image contains rich texture details, including the rocks, the ocean, and the subject’s facial features. Our method closely resembles the Ground Truth, successfully preserving sharp edges and natural colors in the background. H.264, on the other hand, introduces visible artifacts and blurring, particularly around the edges of the subject and the textures of the rocks. The superior detail retention of our approach highlights its capability in handling complex natural scenes with high-frequency textures.
(b) Scene: Person Fishing on Water
This case involves dynamic content, including water reflections and motion. The Ground Truth shows clear water ripples and fine details in the fishing gear and clothing textures. Our method again demonstrates superior visual quality, preserving both the clothing structure and the rippling water with minimal degradation. In contrast, H.264 results in a smeared appearance, especially in the water region, where fine-grained textures are lost. This demonstrates our model’s robustness in maintaining fidelity in motion-heavy, fine-detailed scenes.
(c) Scene: Indoor Interview
This indoor scenario introduces artificial lighting and facial features in a human-centric conversation setup. The Ground Truth image preserves skin tone consistency and clothing folds. Our method produces sharp and visually appealing results that retain fine facial details and contrast. H.264 fails to reproduce the same level of clarity, with noticeable blur in facial features and clothing edges leading to a loss of expression and scene semantics. The results emphasize our model’s advantage in retaining critical human visual cues at ultra low bitrates, even under indoor lighting.
(d) Scene: TEDx Presentation
This example includes stage lighting, text (TEDx letters), and human presence. The Ground Truth shows clear red letters and consistent skin tone under spotlight. Our method maintains high fidelity in both the speaker’s outline and the textured stage background, attribute to segmentation sequences and human motion. The H.264 result suffers from color bleeding and edge artifacts, especially around the large red letters and the person’s silhouette. This shows our model’s capacity to preserve both foreground and background details with high structural accuracy, even under challenging lighting conditions.
In summary, across all four scenes, our method consistently outperforms H.264 by delivering higher perceptual quality and better semantic integrity at ultra low bitrates, which confirms that our method can provide high-quality compressed outputs suitable for applications with heavy transmission constraints, such as satellite video conferencing, surveillance, and media streaming.
@@ Conclusion
This paper presents the first end-to-end video compression framework leveraging conditional diffusion models to achieve human-perception-aligned reconstruction at ultra-low bitrates. Despite these advances, limitations remain: The current decoding speed falls short of real-time requirements due to computational intensity in diffusion-based generation. To address this, we propose adopting familiar model—deploying homologous diffusion models of varying capacities—enabling automatic model switching based on decoder-side computational resources, inspired by AI Flow paradigms (An et al., 2025). Additionally, content-agnostic bitrate control via manual parameter tuning (e.g., $l$, NN, $\xi$) limits adaptability. Future work will integrate content-understanding modules to dynamically optimize compression thresholds based on scene complexity and motion characteristics.

### v4

@@ 3. Conditional Frame Generation at Decoder
Fig. 6 provides a qualitative comparison of video compression performance across different methods, specifically between Ours, H.264/H.265, and the Ground Truth. The selected sample illustrates that our method preserves semantic and structural details even at ultra low bitrates, whereas H.264 and H.265 exhibit significant compression artifacts including noticeable blurring effects.
@@ Conclusion
This paper presents a video compression framework leveraging conditional diffusion models to achieve human-perception-aligned reconstruction at ultra-low bitrates. Despite these advances, limitations remain: The current decoding speed falls short of real-time requirements due to computational intensity in diffusion-based generation. To address this, we propose adopting familiar model—deploying homologous diffusion models of varying capacities—enabling automatic model switching based on decoder-side computational resources, inspired by AI Flow paradigms (An et al., 2025). Additionally, content-agnostic bitrate control via manual parameter tuning (e.g., $l$, NN, $\xi$) limits adaptability. Future work will integrate content-understanding modules to dynamically optimize compression thresholds based on scene complexity and motion characteristics.